At the end of 2017, Ayayai Corporation owns a licence with a remaining life of 11 years and a carrying amount of $505,000. Ayayai expects undiscounted future cash flows from this licence to total $509,900. The licence's fair value is $400,600 and disposal costs are estimated to be nil. The licence's discounted cash flows (that is, value in use) are estimated to be $453,100. Ayayai prepares financial statements in accordance with ASPE. Determine if the licence is impaired at the end of 2017 The licence at the end of 2017. Prepare any related entry that is necessary. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.) Account Titles and Explanation Debit Credit SHOW LIST OF ACCOUNTS Assume the recoverable amount under ASPE (undiscounted future cash flows) is calculated to be $500,000 at the end of 2018. Determine if the licence is impaired at the end of 2018 The licence at the end of 2018.
Prepare any related entry that is necessary. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.) Account Titles and Explanation Debit Credit
Answer #1
1 Carrying value 505000
Fair value 400600
The fair value is less than the carrying amount. It implies that the goodwill is impaired.
Impairment loss is excess of carrying value over recoverable amount
Recoverable amount (Higher of fair value and value in use) 453100
Impairment cost = 505000 - 453100 = 51900
Entry
Particulars Debit ($) Credit ($)
Loss on impairment 51900
Licence 51900
To record loss in impairment of licence
2 Recoverable amount 500000
Impairment loss = 505000 - 500000 = 5000
Entry
Particulars Debit ($) Credit ($)
Loss on impairment 5000
Licence 5000
To record loss in impairment of licence